OpenStudy (anonymous):

5x + 2(x - 3) = -2(x - 1)

OpenStudy (anonymous):

5x + 2(x - 3) = -2(x - 1) Distribute the 2 over to (x -3) and the -2 over to (x-1)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

wwhat now

OpenStudy (anonymous):

Are you trying to simplify this?

OpenStudy (anonymous):

5x+2x-6=-2x+2 7x-6=2x+2 7x=2x+8 5x=8 x=8/5 If you are solving for x